OpenStudy (anonymous):

Simplify: 2-2 tan (theta) cot(theta)/ (1-tan(theta)) (1-cot(theta))

OpenStudy (anonymous):

Is the equation \[{2-2tan(\theta)cot(\theta)\over (1-tan(\theta)(1-cot(\theta)}\]?

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (anonymous):

If that is the case, then it simplifies to zero because \[cot(\theta) = {1\over tan(\theta)}\]so the numerator becomes\[2-2{tan(\theta)\over tan(\theta)}\]which is zero

OpenStudy (anonymous):

Ok, that was the answer I came up with but I just wanted to make sure i did it right.

OpenStudy (anonymous):

Yep, you sure did

OpenStudy (anonymous):

Thanks so much!